# Seeded Topology Optimization for Commercial Foundry Integrated Photonics

### v1

@@ 5 Conclusion
We have demonstrated a seeded TO design methodology that allows for optimization of integrated photonic devices that yields improved performance over traditional TO. The seed is best chosen as a known functional device created via traditional TO, other optimization methods such as parameter or shape optimization, or from physics based models. Seeded TO relies on a blurring filter chosen to perturb the known structure seeking an optimized design. This new design methodology enables the greation of new, more robust algorithms to ensure the device meets DRC. We illustrated this optimization technique using four different test devices designed for a foundry process.
@@
Seeded TO brings several important benefits to the integrated photonics inverse design community not seen in traditional topology optimization implementations. While traditional TO has the capability to discover non-intuitive device geometries, the optimal topology changes throughout the optimization as the design parameters binarize and DRC constraints are applied. This limits the ability of TO to find a strong optimal topology without careful hyperparameter tuning or a post-TO optimization scheme. Less hyperparameter tuning is required when using seeded TO resulting in fewer reloads throughout the optimization and a reduced total computational cost. This allows the designer to focus on the other aspects of the design process as hyperparameter tuning is often a tedious, time-consuming process. Seeded TO is compatible with any features implemented in TO including multi-layer and non-vertical sidewall optimization [35, 20]. A desirable consequence of seeded TO is the elimination of the periodic ripples that are commonly develop through traditional TO and have limited impact on device performance. This reduces the effect lithographic smoothing has on the device topology, resulting in greater alignment between measured and simulated performance.
@@
There remain many opportunities for future work including exploring potential design techniques that can be used to create the seed (e.g. inverse design, shape optimization, physics-defined design, etc.). Designing a seed using using a physics-defined topology will allow for the optimization of large structures such as multimode interferometers, spot-size converters, and Bragg grating filters that have traditionally been difficult to inverse design due to simulation complexity. Additionally, there is scope to explore additional functionality in seeded TO such as creating algorithms that identify non-essential features of TO structures that can be removed to reduce device footprint. Intelligent implementation of non-gradient based design parameter modification has scope to significantly improve the performance of devices designed using TO.
Appendix A: High Contrast Grating Reflectance From Resonator Response
Figure 8: (a) Add/drop ring resonator and (b) high contrast grating (HCG) resonator with input, through, and drop ports labeled.
@@
where $\kappa=\sqrt{1-t^{2}}$ is the directional coupler cross-coupling coefficient, $t$ is the directional coupler self-coupling coefficient (assumed real), $A$ is the round-trip optical power attenuation, and $\phi$ is the round-trip optical phase [32]. The transmission to the drop port is:
$T_{d}\left(\phi\right)=\frac{I_{drop}}{I_{in}}=\left|\frac{E_{drop}}{E_{in}}\right|^{2}=\frac{\kappa^{4}\sqrt{A}}{1-t^{2}\sqrt{A}\left(e^{j\phi}+e^{-j\phi}\right)+t^{4}A}=\frac{\kappa^{4}\sqrt{A}}{1-2t^{2}\sqrt{A}\cos\phi+t^{4}A}$ (13)
@@
$T_{d}\left(\phi\right)=\frac{T_{d,max}\left(1-Rt^{2}\right)^{2}}{1+R^{2}t^{4}-2Rt^{2}(1-2\sin^{2}\left(\phi/2\right))}=\frac{T_{d,max}\left(1-Rt^{2}\right)^{2}}{\left(1-Rt^{2}\right)^{2}+4Rt^{2}\sin^{2}\left(\phi/2\right)}$ (17)
@@
$T_{d}\left(\Delta\phi_{FWHM,d}/2\right)=\frac{T_{d,max}}{2}\rightarrow\frac{T_{d,max}}{1+F\sin^{2}\left(\Delta\phi_{FWHM,d}/4\right)}=\frac{T_{d,max}}{2}$ (20)
@@
The phase bandwidth is related to the FWHM ($\Delta\lambda_{FWHM}$) and the FSR ($\Delta\lambda_{FSR}$) by [41]:
@@
[…]
Funding This material is based upon work supported in part by the National Science Foundation (NSF) Center “EPICA” under Grant No.1 2052808, https://epica.research.gatech.edu/. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the NSF. JMH, CAK, JJW, PA, and SER were supported by the Georgia Electronic Design Center of the Georgia Institute of Technology.
[…]
Acknowledgments This research was supported in part through research cyberinfrastructure resources and services provided by the Partnership for an Advanced Computing Environment (PACE) at the Georgia Institute of Technology, Atlanta, Georgia, USA [42]. The authors would like to thank GlobalFoundries for providing silicon fabrication through GF45CLO university program.
[…]
Disclosures The authors declare no conflicts of interest.
[…]
Data availability Data underlying the results presented in this paper are not publicly available at this time but may be obtained from the authors upon reasonable request.

### v5

@@ V Conclusion
We have demonstrated a seeded TO design methodology that enables effective optimization of commercial foundry-compatible integrated photonic devices, yielding improved performance over traditional TO. The seed is best chosen as a known functional device created via traditional TO, other optimization methods such as parameter or shape optimization, or from physics-based models. Seeded TO relies on a blurring filter chosen to perturb the known structure seeking an optimized design. This new design methodology enables the creation of new, more robust algorithms to ensure the device meets DRC. We illustrated this optimization technique using four different test devices designed for a foundry process.
@@
Seeded TO brings several important benefits to the integrated photonics inverse design community not seen in traditional topology optimization implementations. While traditional TO has the capability to discover non-intuitive device geometries, the optimal topology changes throughout the optimization as the design parameters binarize and DRC constraints are applied. This limits the ability of traditional TO to find a strong optimal topology without careful hyperparameter tuning or a post-TO optimization scheme. In many TO implementations hyperparameter tuning is done heuristically by executing a TO algorithm and evaluating the solution upon completion Ha and Carstensen (2024); Lynch, Sarkar, and Maute (2019). Less hyperparameter tuning is required when using seeded TO resulting in fewer reloads throughout the optimization, potentially reducing the computational cost of an optimization. This allows the designer to focus on the other aspects of the design process as hyperparameter tuning is often a tedious, time-consuming process. The computational cost of an optimization can be further reduced by performing traditional TO using low-resolution or 2D FDTD simulations followed by full resolution seeded TO Hiesener et al. (2025).
Though only DRC fabrication constraints were considered in this work, additional TO constraints and permittivity mapping schemes are compatible with seeded TO. This includes constraints on the etching process for multilayer designs and permittivity projection operations used to optimize devices on platforms with non-vertical sidewalls Hammond et al. (2022b); Probst et al. (2024); Shang et al. (2023); Probst et al. (2025); Pan et al. (2021). A desirable consequence of seeded TO is the elimination of the periodic ripples that are commonly develop through traditional TO and have limited impact on device performance. This reduces the effect lithographic smoothing has on the device topology, resulting in greater alignment between measured and simulated performance.
@@
There remain many opportunities for future work including exploring potential design techniques that can be used to create the seed (e.g. inverse design, shape optimization, physics-defined design, etc.). Designing a seed using a physics-defined topology will allow for the optimization of large structures such as multimode interferometers, spot-size converters, and Bragg grating filters that have traditionally been difficult to inverse design due to simulation complexity. Microcavity design using local density of states is a common design problem in TO which is sensitive to small perturbations and can develop tiny features that prevent fabrication Liang and Johnson (2013), a seeded TO method may assist in improving the fabricability and robustness of these devices. Since seeded TO performs filtering and DRC correction outside of TO, additional constraints such as requiring all features of a device to be connected can easily be incorporated into seeded TO Li et al. (2016); Christiansen and Sigmund (2021). There is scope to explore additional functionality in seeded TO such as creating algorithms that identify non-essential features of TO structures that can be removed to reduce device footprint. Intelligent implementation of non-gradient based design parameter modification has scope to significantly improve the performance of devices designed using TO.
VI Supplemental Material
VI.1 TO Optimizers
TO typically requires a gradient-based nonlinear optimization method to calculate the design parameter step for each iteration. Two commonly employed methods are the method of moving asymptotes (MMA) or L-BFGS-B Svanberg (2002); Zhu et al. (1997). In this work we employ the globally convergent method of moving asymptotes; however, different optimizers have different constraints and produce different optimization trajectories which need to be considered when designing for a specific application.
MMA uses simple functions with penalty terms to approximate the nonlinear function and determine the next step. L-BFGS-B uses the gradient of the objective function and a limited-memory approximation of the hessian derived from previous iterations to determine the descent direction in the presence of bound constraints. Both algorithms are well-suited for large-scale optimization problems with thousands of design variables. However, L-BFGS-B only accepts bound constraints i.e. constraints where the design variable falls between scalar limits. MMA, on the other hand, also permits constraints that are nonlinear functions of the design variables, allowing more freedom in choice of constraints. The commercial foundry integrated photonics design problem introduces many foundry-based design constraints an MMA optimizer can handle well.
L-BFGS-B applied to photonics problems tends to produce optimization paths that are very nearly monotonic since it exploits Hessian information to obtain an accurate update direction Vercruysse et al. (2019). Furthermore, due to the update precision L-BFGS-B may require only 100 iterations to stop improving the objective function significantly. MMA takes suboptimal steps that result in optimization paths with spikes Ballew et al. (2023); Hammond et al. (2021). The spikes depend greatly on the specific problem being solved, for example bends contain few while a broadband mirror has many Hammond et al. (2021). MMA also requires on the order of 200 iterations to produce an optimized design. There are many other optimizers explored in the inverse design community, each with benefits and drawbacks that make no optimizer ideal for all integrated photonic applications.
VI.2 Foundry-Set Constraint Implementations
In this section we outline our design rule check (DRC) constraint implementation for geometric linewidth constraints (GLC) and area constraints (AC).
VI.2.1 Geometric Linewidth Constraints
GLC includes the minimum linewidth and linespacing which are the minimum lengthscales of solid and void features that can be accurately fabricated. These constraints are commonly combined and referred to as the minimum feature size; however, many platforms have different values for the minimum linewidth and linespacing, making it useful to separate these constraints. The minimum linewidth constraint ($g_{LW}$) is described by the function
$g_{LW}=\frac{1}{n}\sum_{i\in N}I_{i}^{WL}(\rho_{i})\cdot\left[\text{min}\{\left(\tilde{\rho}-\eta_{e}\right),0\}\right]^{2}$ (16)
where $n$ is the number of inflection regions ($N$) identified that violate the minimum linewidth, $\bar{\rho}$ is the projected design parameters, $I_{i}^{WL}(\rho_{i})$ is the indicator function that identifies each inflection region of the solid phase, and $\eta_{e}$ is the linewidth threshold parameter Hammond et al. (2021). The indicator function is defined as
$I_{i}^{LW}(\rho)=\bar{\rho}\cdot\exp{\left(-c\left|\nabla\tilde{\rho}\right|^{2}\right)}$ (17)
where $c$ is a dampening term that dictates the strength of the indicator function, this is typically set to $r^{4}$ where $r$ is the design grid resolution Hammond et al. (2021). The linewidth threshold parameter is given by
$\eta_{e}=\begin{cases}\frac{1}{4}\left(\frac{l_{W}}{R}\right)^{2}+\frac{1}{2}&\frac{l_{W}}{R}\in\left[0,1\right]\\ -\frac{1}{4}\left(\frac{l_{W}}{R}\right)^{2}+\frac{l_{W}}{R}&\frac{l_{W}}{R}\in\left[1,2\right]\\ 1&\frac{l_{W}}{R}\in\left[2,\infty\right]\end{cases}$ (18)
where $l_{w}$ is the minimum linewidth and $R$ is the user-specified radius of the conic filter Hammond et al. (2021). This allows the user to arbitrarily choose the filter radius without dependence on the foundry constraints. Similarly, the minimum linespacing constraint ($g_{LS}$) is described by
$g_{LS}=\frac{1}{n}\sum_{i\in N}I_{i}^{LS}(\rho_{i})\cdot\left[\text{min}\{\left(\eta_{d}-\tilde{\rho}\right),0\}\right]^{2}$ (19)
where the indicator function ($I_{i}^{LS}(\rho_{i})$) which identifies the inflection region of the void phase is given by
$I_{i}^{LS}(\rho)=\left(1-\bar{\rho}\right)\cdot\exp{\left(-c\left|\nabla\tilde{\rho}\right|^{2}\right)}$ (20)
The linespacing threshold parameter ($\eta_{d}$) is
$\eta_{d}=\begin{cases}\frac{1}{2}-\frac{1}{4}\left(\frac{l_{S}}{R}\right)^{2}&\frac{l_{S}}{R}\in\left[0,1\right]\\ 1+\frac{1}{4}\left(\frac{l_{S}}{R}\right)^{2}-\frac{l_{S}}{R}&\frac{l_{S}}{R}\in\left[1,2\right]\\ 0&\frac{l_{S}}{R}\in\left[2,\infty\right]\end{cases}$ (21)
where $l_{S}$ is the minimum linespacing Qian and Sigmund (2013). Using circular filters, the linewidth and linespacing implementations both impose implicit constraints on the corresponding minimum curvature for both solid and void regions. The minimum radius of curvature is given by Hammond et al. (2021)
$k_{W,S}=\frac{l_{W,S}}{2}$ (22)
VI.2.2 Minimum Area Constraints
AC includes the minimum area and enclosed area constraints, which dictate the smallest allowable island or hole that can be accurately fabricated. The minimum area constraint function ($g_{A}$) is defined as
$g_{A}=\sum_{i\in N}\sin\left(\frac{\pi}{A_{min}}A_{i}\left(\bar{\rho},I_{A}(\bar{\rho})\right)\right)$ (23)
where $N$ are the contours of the topology that violate the minimum area, $A_{min}$ is the minimum area, $A_{i}\left(\bar{\rho},I_{A}(\bar{\rho})\right)$ is the area of the $i^{th}$ contour, $\bar{\rho}$ are the projected design parameters and $I_{A}(\bar{\rho})$ is an indicator function that is marks all regions of the topology that contain islands with areas below the minimum area constraint. The minimum enclosed area constraint function ($g_{EA}$) is defined as
$g_{EA}=\sum_{i\in M}\sin\left(\frac{\pi}{E_{min}}E_{i}\left(1-\bar{\rho},I_{EA}(1-\bar{\rho})\right)\right)$ (24)
where $M$ is the contours of the topology that violate the minimum enclosed area, $E_{min}$ is the minimum enclosed area, $E_{i}\left(1-\bar{\rho},I_{EA}(1-\bar{\rho})\right)$ is the area of the $i^{th}$ contour, and $I_{EA}(1-\bar{\rho})$ is an indicator function that marks all regions of the topology that contain holes with areas below the minimum enclosed area constraint.
The indicators ($I_{A}$, $I_{EA}$) are determined using an out-of-the box python package to perform the marching-squares algorithm to extract contours from the design parameters ($\bar{\rho}$, $1-\bar{\rho}$) Van der Walt et al. (2014); Lorensen and Cline (1998). With the contours identified, the area of each contour ($A_{i}$, $E_{i}$) is calculated using another out-of-the box Python function that calculates areas using a discrete summation of all the density values inside the contour, which are identified using morphological dilations Virtanen et al. (2020); Hammond et al. (2021). If a particular contour has a smaller area than the minimum area, the filled contour region is dilated by 1 pixel and added to the indicator function.
The constraints are defined as
$g_{A,EA}\leq 0$ (25)
such that the optimizer drives each constraint to 0 Hammond et al. (2021). The $\sin$ in the constraint functions enables both erosion and dilation of islands and holes depending on the size of the violating feature. If the area of an island is less than half of the minimum area, it is eroded, whereas if the area is larger than half the minimum area, it is expanded.
VI.3 High Contrast Grating Reflectance From Resonator Response
@@
Figure 8: (a) Add/drop ring resonator and (b) high contrast grating (HCG) resonator with directional couplers (DC), input ports, through ports, and drop ports labeled.
where $\kappa=\sqrt{1-t^{2}}$ is the directional coupler cross-coupling coefficient, $t$ is the directional coupler self-coupling coefficient (assumed real), $A$ is the round-trip optical power attenuation, and $\phi$ is the round-trip optical phase Chrostowski and Hochberg (2015). The transmission to the drop port is:
$T_{d}\left(\phi\right)=\frac{I_{drop}}{I_{in}}=\left|\frac{E_{drop}}{E_{in}}\right|^{2}=\frac{\kappa^{4}\sqrt{A}}{1-t^{2}\sqrt{A}\left(e^{j\phi}+e^{-j\phi}\right)+t^{4}A}\\ =\frac{\kappa^{4}\sqrt{A}}{1-2t^{2}\sqrt{A}\cos\phi+t^{4}A}$ (27)
@@
$T_{d}\left(\phi\right)=\frac{T_{d,max}\left(1-Rt^{2}\right)^{2}}{1+R^{2}t^{4}-2Rt^{2}(1-2\sin^{2}\left(\phi/2\right))}=\\ \frac{T_{d,max}\left(1-Rt^{2}\right)^{2}}{\left(1-Rt^{2}\right)^{2}+4Rt^{2}\sin^{2}\left(\phi/2\right)}$ (31)
@@
$T_{d}\left(\Delta\phi_{FWHM,d}/2\right)=\frac{T_{d,max}}{2}\rightarrow\\ \frac{T_{d,max}}{1+F\sin^{2}\left(\Delta\phi_{FWHM,d}/4\right)}=\frac{T_{d,max}}{2}$ (34)
@@
The phase bandwidth is related to the FWHM ($\Delta\lambda_{FWHM}$) and the FSR ($\Delta\lambda_{FSR}$) by Van (2016):
@@
Acknowledgements.
This material is based upon work supported in part by the National Science Foundation (NSF) Center “EPICA” under Grant No.1 2052808, https://epica.research.gatech.edu/. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the NSF. JMH, CAK, JJW, PA, and SER were supported by the Georgia Electronic Design Center of the Georgia Institute of Technology. This research was supported in part through research cyberinfrastructure resources and services provided by the Partnership for an Advanced Computing Environment (PACE) at the Georgia Institute of Technology, Atlanta, Georgia, USA PACE (2017). The authors would like to thank GlobalFoundries for providing silicon fabrication through FotonixTM university program.
Data Availability Statement
The data that support the findings of this study are available from the corresponding author upon reasonable request.